# Self-supervised Brain Lesion Generation for Effective Data Augmentation of Medical Images

### v1

@@ 4.3.2 Downstream Segmentation Model Performance
We compare the segmentation model performance using our framework to other SOTA segmentation models. Note that our framework can use any backbone model to perform the segmentation task, here we consider both UNet and Attention UNet as the base segmentation models. Segmentation performance for all models was evaluated using the Dice similarity coefficient (DSC), average surface distance (ASD), and 95% Hausdorff distance (HD95). Table 4 shows the segmentation model performance for the ATLAS v2.0 dataset. Our approach consistently demonstrates improved performance for all metrics compared to all competing models, irrespective of the base model. This consistent overperformance is attributed to two pivotal enhancements: the use of an augmented brain lesion dataset, which closely mimics real-world appearance for model training, and the incorporation of the prototype consistency loss to align model features for real and synthetic lesions. These enhancements not only elevate the base model’s ability to accurately segment images but also emphasize a crucial insight: the quality of the dataset plays a more critical role than the network architecture in achieving model generalization for segmentation tasks. Our approach, by leveraging high-fidelity synthetic data and strategic feature alignment, effectively unleashes the potential of classical segmentation models like UNet, establishing that dataset augmentation and targeted modifications to the training loss function can improve SOTA models for brain stroke lesion segmentation.
@@
Table 4: Quantitative results of our method and other segmentation models for the ATLAS v2.0 dataset. * indicates that the p-value ¡ 0.05 compared to the second-best model using a paired Student’s t-test.
@@
Ours (UNet) $\textbf{60.23$\pm$29.48}^{*}$ $\textbf{6.32$\pm$13.68}^{*}$ $\textbf{20.26$\pm$25.81}^{*}$
Ours (Attention UNet) 57.69$\pm$29.29 7.53$\pm$15.60 21.35$\pm$27.27
The quantitative metrics for model segmentation performance in the Shift MS Dataset are reported in Table 3. As with ATLAS v2.0, our framework has improved performance compared to the other models for the out-domain development set (Devout), which demonstrates our framework improves domain generalization.
@@
Figure 5: Visualization of lesion segmentation in the ATLAS v2.0 dataset for different models. Here ’Att UNet’ is short for the Attention UNet backbone.
Figure 6: Visualization of lesion segmentation on the Shift MS dataset for different models. Here ’Att UNet’ is short for the Attention UNet backbone.
@@ 4.3.3 Comparisons with Different Data Augmentation Methods
We compare the performance of our framework with other data augmentation methods, including voxel-based methods and GAN-based methods. Voxel-based data augmentation methods use a combination of two existing images to create a new synthetic image. Here we adopted six methods: i.e. Mixup [3], CutMix [4], Copy-Paste [14], TumorCP [41], CarveMix [3], and SelfMix [42]. For the GAN-based methods [43], a conditional GAN was adapted to generate either a deformation field (D), intensity field (I), or a combination of both (D+I) to change the structure and/or the appearance of images. Regardless of the data augmentation method, UNet was the segmentation model used.
@@
Table 5: UNet Segmentation model performance for our method and other data augmentation techniques on the ATLAS v2.0 dataset. For the GAN method, we consider deformation augmentation (D), intensity augmentation (I), and both (D+I). * indicates that our framework significantly improves segmentation performance (p-value ¡ 0.05) compared to the second-best model using a paired Student’s t-test.
@@
Table 6: UNet Segmentation model performance for our framework and other data augmentation techniques on the Shift MS dataset. For the GAN method, we consider deformation augmentation (D), intensity augmentation (I), and both (D+I). * indicates that our framework significantly improves segmentation performance (p-value ¡ 0.05) compared to the second-best model using a paired Student’s t-test.
@@ 4.4 Ablation Studies
We validate the effectiveness of the three modules in our framework i.e. Lesion synthesis(SL), Soft Poisson Blending (SPB), and Prototype Consistency (PC). All ablation studies were conducted on the ATLAS v2.0 dataset and Shift MS dataset.
@@ 4.4.3 Prototype Consistency
5 Conclusions
We presented a comprehensive framework to augment existing training samples for brain lesion segmentation via an efficient generative model, Soft Poisson Blending (SPB) to create a seamless boundary between foreground and background, and finally a prototype consistency regularisation term to ensure the segmentation model learns similar features for synthetic and real lesions. The generative model is trained in a self-supervised manner but can generate synthetic lesions with the same latent space distribution as real lesions. SPB eliminates boundary artifacts when inserting synthetic lesions into existing brain images to create fake yet realistic training samples. The augmented samples boost segmentation model performance under the supervision of the prototype consistency penalty. The penalty helps to learn the aligned features of real and synthetic lesions to achieve a better segmentation performance compared to other models and data augmentation approaches. Experiments on two public datasets demonstrate that our framework outperforms other methods that only adapt augmented samples for model training. Currently, our framework is only validated on brain lesion MRI datasets. Extending our framework to other image modalities and other organs will be future work. Additionally, we will explore replacing the GAN-based lesion synthesis module with a diffusion model to create even more realistic images for data augmentation in the future.

### v2

@@ 4.3.2 Downstream Segmentation Model Performance
We compare the segmentation model performance using our framework to existing segmentation models. All models were trained from scratch. Note that our framework can use any backbone model to perform the segmentation task, here we consider UNet as the base segmentation model. Segmentation performance for all models was evaluated using the Dice similarity coefficient (DSC), average surface distance (ASD), and 95% Hausdorff distance (HD95). Table 4 shows the segmentation model performance for the ATLAS v2.0 dataset. Our approach consistently demonstrates improved performance for all metrics compared to all competing models, irrespective of the base model. This consistent overperformance is attributed to two pivotal enhancements: the use of an augmented brain lesion dataset, which closely mimics real-world appearance for model training, and the incorporation of the prototype consistency loss to align model features for real and synthetic lesions. These enhancements not only elevate the base model’s ability to accurately segment images but also emphasize a crucial insight: the quality of the dataset plays a more critical role than the network architecture in achieving model generalization for segmentation tasks. Our approach, by leveraging high-fidelity synthetic data and strategic feature alignment, effectively unleashes the potential of classical segmentation models like UNet, establishing that dataset augmentation and targeted modifications to the training loss function can improve segmentation models’ performances for brain stroke lesion segmentation.
@@
Table 4: Quantitative results of our method and other segmentation models for the ATLAS v2.0 dataset. * indicates that the p-value $<$ 0.05 compared to the second-best model using a paired Student’s t-test.
@@
Ours $\textbf{60.23$\pm$29.48}^{*}$ $\textbf{6.32$\pm$13.68}^{*}$ $\textbf{20.26$\pm$25.81}^{*}$
The quantitative metrics for model segmentation performance in the Shift MS Dataset are reported in Table 3. As with ATLAS v2.0, our framework has improved performance compared to the other models for the out-domain development set (Devout), which demonstrates our framework improves domain generalization ability.
@@
Fig. 6: Visualization of lesion segmentation in the ATLAS v2.0 dataset for different models. Here ’Att UNet’ is short for the Attention UNet.
@@ 4.3.3 Comparisons with Different Data Augmentation Methods
We compare the performance of our framework with other data augmentation methods, including voxel-based methods and GAN-based methods. Voxel-based data augmentation methods use a combination of two existing images to create a new synthetic image. Here we adopted six methods: i.e., Mixup [43], CutMix [42], Copy-Paste [6], TumorCP [41], CarveMix [43], and SelfMix [45]. For the GAN-based methods [4], a conditional GAN was adapted to generate either a deformation field (D), intensity field (I), or a combination of both (D+I) to change the structure and/or the appearance of images. Regardless of the data augmentation method, UNet was the segmentation model used.
@@
Fig. 7: Visualization of lesion segmentation on the Shift MS dataset for different models. Here ’Att UNet’ is short for the Attention UNet.
Table 5: UNet Segmentation model performance for our method and other data augmentation techniques on the ATLAS v2.0 dataset. For the GAN method, we consider deformation augmentation (D), intensity augmentation (I), and both (D+I). * indicates that our framework significantly improves segmentation performance (p-value $<$ 0.05) compared to the second-best model using a paired Student’s t-test.
@@
Table 6: UNet Segmentation model performance for our framework and other data augmentation techniques on the Shift MS dataset. For the GAN method, we consider deformation augmentation (D), intensity augmentation (I), and both (D+I). * indicates that our framework significantly improves segmentation performance (p-value $<$ 0.05) compared to the second-best model using a paired Student’s t-test.
@@ 4.4 Ablation Studies
We validate the effectiveness of the three modules in our framework i.e., Lesion synthesis (SL), Soft Poisson Blending (SPB), and Prototype Consistency (PC). All ablation studies were conducted on the ATLAS v2.0 dataset and Shift MS dataset.
@@ 4.4.3 Prototype Consistency
5 Conclusions & Discussions
We presented a comprehensive framework to augment existing training samples for brain lesion segmentation via a two-stage adversarial autoencoder (AAE) to generate new lesion images. The AAE is trained in a self-supervised manner, but generates synthetic lesions with the same latent space distribution as real lesions. We then augment the synthetic images by using Soft Poisson Blending (SPB) to create a seamless boundary between foreground and background, eliminating boundary artifacts. Finally we introduce a prototype consistency regularisation term during segmentation model training to ensure similar features across synthetic and real lesions are learnt. The synthetic lesion samples boost segmentation model performance under the supervision of the prototype consistency penalty. Experiments on two public datasets demonstrate that our framework outperforms other data augmentation approaches and methods that only adapt augmented samples for model training. We do not compare our approach to models based on pre-trained datasets such as SAM [14] because they are pre-trained on a large-scale dataset, making direct comparisons unfair. Besides, SAM-based methods largely depend on accurate user prompts to achieve good segmentation results, which differs from our fully automatic setting requiring no prompt. Currently, our framework is validated on brain lesion MRI datasets. Extending our framework to other image modalities and other organs will be future work. Additionally, we will explore adding conditions to further control the process of lesion image synthesis for controllable data augmentation in the future.
Acknowledgments
This work was supported by Centre for Doctoral Training in Surgical and Interventional Engineering at King’s College London; the funding from the Wellcome Trust Award (218380/Z/19/Z) and the Wellcome/EPSRC Centre for Medical Engineering (WT203148/Z/16/Z).